Calc. the pH of 10^-8 M acid sol.

Grade:12

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Find the dissociation constant first and then write log[H+]

pH=-log[H+]

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

10-8M acid = 10-8M H+ ion concentration...

in solution H+ from water is 10-7 M

total H+ = 10-7+10-8 = 1.01*10-7

pH = -log[H+] =-log[1.01*10-7]

    = 6.85       approx

Sudheesh Singanamalla
13 years ago

Molarity = 10^-8 .

n/V = 10^-8

n = 10^-8 * V

pH = -log [H+] .

Water normally dissociates: H2O <===> H+ + OH-

The dissociation constant is Kw = 10^-7

So you would have 10^-7 moles of H+ in pure water. Add the 10^-8 moles of acid, and the resulting concentration of H+ is [H+] = 1.1 x 10^-7

the negative log of 1.1 x 10^-7 = 6.96, which is the pH value.

therefore pH of the solution is 6.96

----------------------------------------------------------------------------------------------------------------------

PLEASE APPROVE

----------------------------------------------------------------------------------------------------------------------